OpenStudy (anonymous):

these are hard to put on here but what is the solution to the system of equations |{ -3x-y+2z=-4 <|{ x-y+z=10 |{ 2x-3y+z=4

OpenStudy (amistre64):

matrix or sub seems to be the easiest.... sub being the eaiser

OpenStudy (amistre64):

-3x-y+2z=-4 x-y+z=10 <-- z = 10-x+y 2x-3y+z=4 -3x -y +2(10-x+y) = -4 -3x -y +20 -2x+2y = -4 -5x +y = -24 ; y = 5x -24 ----------------------- 2x -3y+(10-x+y) = 4 x -2y = -6 ------------------------ x - 2(5x -24) = -6 x - 10x +48 = -6 -9x = -54 x = 6 ................................. y = 5x -24 y = 30 -24 = 6 ................................. z = 10-x+y z = 10-6+6 = 10 dbl check

OpenStudy (anonymous):

i am

OpenStudy (amistre64):

-3x-y+2z=-4 -3(6) -6 +2(10) = -4 -18 -6 +20 = -4 -24 +20 = -4 -4 = -4 ...checks x-y+z=10 6-6+10 = 10 0+10=10 10=10 .... checks 2x-3y+z=4 2(6) -3(6) +10 = 4 12 -18 +10 = 4 -6 +10 = 4 4=4 ......checks

OpenStudy (amistre64):

x=6 , y=6 , z=10